2 charges eachof q=+2 micro coloumb are placed at the vertices Q and R of the triangle .the sum of sides PQ and PR is 12cm and their product is 32 cm_2 .the potential at point P is

Grade:12th Pass

1 Answers

Yogita Bang
11 years ago

In ΔPQR, Let PQ=8cm PR=4cm

At P, V=9*109 {(2/4) + (2/8)}*10-6*102 = 9*109*(3/4)*10-4 = 6.75*105

 

 

If u like the answer CLICK YES below and approve